OpenStudy (anonymous):

equation: e^2x+5 = 8

OpenStudy (anonymous):

e^2x+5 = 8 e^2x=3 ln(e^2x)=ln3 2x=ln3 x=(ln3)/2=0.5493

OpenStudy (anonymous):

It is e^2x+5 = 8 not e^(2x+5) = 8 right?

OpenStudy (anonymous):

yhea like you first said